### v1

@@ 4 Equivalence of the Aubin property and the strong regularity
This section establishes the main result of this paper. Let ${\boldsymbol{x}}^{*}$ be a locally optimal solution to (1.1) with ${\boldsymbol{y}}^{*}=({\bm{\lambda}}^{*},{\bm{\mu}}^{*})$ being the associated multiplier. Note that any index $j\in\{1,\ldots,J\}$ belongs to one and only one of the following $6$ sets:
@@
For simplicity, $J_{1}\cup J_{2}$ is abbreviated to $J_{1,2}$, $J_{1}\cup J_{2}\cup J_{3}$ is abbreviated to $J_{1,2,3}$, and etc. For a index set $I\subseteq\{1,2,3,4,5,6\}$, $|J_{I}|$ denotes the cardinality of the set $J_{I}$. Note that when $r_{j}=0$, $j$ must belong to $J_{1,2,6}$. Recall from [5, Lemma 25] that, for each $j=1,\ldots,J$, the tangent cone of ${\mathcal{Q}}_{j}$ at $g^{j}({\boldsymbol{x}}^{*})$ is given by
@@
Furthermore, by [5, Corollary 26] the critical cone ${\mathcal{C}}({\boldsymbol{x}}^{*})$ can be given by
@@
Moreover, according to [5, eq. (47)], the affine hull of the critical cone ${\mathcal{C}}({\boldsymbol{x}}^{*})$, i.e. the smallest subspace that contains the critical cone, can be formulated as
@@
The following definition on second-order and strong second-order sufficient conditions for (1.1) comes from [5, Theorems 29 & 30].
@@
Let ${\boldsymbol{x}}^{*}$ be a stationary point of the nonlinear SOCP (1.1). We say that the second-order sufficient condition holds at ${\boldsymbol{x}}^{*}$ if there exists a associated multiplier $({\bm{\lambda}}^{*},{\bm{\mu}}^{*})$ such that
@@
where $\nabla_{{\boldsymbol{x}}{\boldsymbol{x}}}^{2}{\mathcal{L}}$ is the Hessian of ${\mathcal{L}}$ with respect to $x$, and ${\mathcal{H}}({\boldsymbol{x}}^{*},{\bm{\mu}}^{*}):=\sum_{j=1}^{J}{\mathcal{H}}_{j}({\boldsymbol{x}}^{*},{\bm{\mu}}^{*})$ with
@@
We say that the strong second-order sufficient condition holds at ${\boldsymbol{x}}^{*}$ if there exists a associated multiplier $({\bm{\lambda}}^{*},{\bm{\mu}}^{*})$ such that
@@ 4.1 Implications of the Aubin property
The following lemma provides a convenient tool for characterizing the affine hull of the critical cone (4.4) via the null space of a given matrix.
@@
Suppose that Assumption 4.1 holds. Let $\Theta\in{\mathds{R}}^{n\times n}$ be the matrix given in (4.11), and $H\in{\mathds{R}}^{n\times\ell}$ be a matrix with full column rank such that ${\sf rge}H={\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$. Then the mapping ${{\mathbb{S}}}_{\rm GE}$ in (1.8) has the Aubin property at the origin for ${\boldsymbol{x}}^{*}$ implies that for all ${\bm{\nu}}\in{\mathds{R}}^{\ell}\setminus\{{\boldsymbol{0}}\}$ it holds that
@@ 4.2 Deriving the strong second-order sufficient condition
Since the solution mapping ${{\mathbb{S}}}_{\rm KKT}$ in (1.5) has the Aubin property at $({\boldsymbol{0}},{\boldsymbol{0}})$ for $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$, we know that the mapping ${\mathbb{S}}_{\rm GE}$ in (1.8) also has the Aubin property at the origin for ${\boldsymbol{x}}^{*}$. Then, by [8, corollary 25] the constraint nondegeneracy condition (1.7) holds at ${\boldsymbol{x}}^{*}$ and the second-order sufficient condition holds (4.5) at ${\boldsymbol{x}}^{*}$. Therefore, Assumption 4.1 holds.
@@ 4.3 The equivalence theorem
Note that (1) is an immediate consequence of (2). On the other hand, if (1) holds, we know from [8, Corollary 25] that ${\boldsymbol{x}}^{*}$ is nondegenerate and the second-order sufficient condition (4.5) holds at ${\boldsymbol{x}}^{*}$. Consequently, it follows from Proposition 4.3 that the strong second-order sufficient condition (4.6) holds. Thus, by using [5, Theorem 30] we know that (2) holds. This completes the proof. ∎
Theorem 4.1 establishes the equivalence between the Aubin property and the strong regularity for the nonlinear SOCP (1.1) at local optimal solutions without requiring the strict complementarity. Moreover, Theorem 4.1 implies its counterpart for conventional nonlinear programming (((([10, Theorems 1, 4 &\&& 5])))) as a special case as one can take $r_{j}=0$ for all $j=1,\dots,J$.
According to Theorem 4.1, ${{\mathbb{S}}}_{\rm KKT}$ in (1.5) has the Aubin property at $({\boldsymbol{0}},{\boldsymbol{0}})$ for $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$ is also equivalent to many other conditions, such as (a) ${\boldsymbol{x}}^{*}$ is nondegenerate and the uniform second order growth condition ([4, Definition 5.16]) holds at ${\boldsymbol{x}}^{*}$ (by [5, Theorem 30]); (b) ${\boldsymbol{x}}^{*}$ is a nondegenerate and fully stable ([21, Definition 4.1]) locally optimal solution (by [21, Theorem 4.8]); and (c) Clarke’s generalized Jacobian of a KKT system is nonsingular (by [31, Theorem 3.1]). One may refer to the relevant references for details.
@@ 5 Conclusions
In this paper, we have established the equivalence, between the Aubin property of the perturbed KKT system and the strong regularity of the KKT system for nonlinear SOCP problems, at locally optimal solutions. Such a result removes the restriction that $|J_{4}\cup J_{5}\cup J_{6}|\leq 1$ in [23] and [22]. We achieved this by introducing a novel reduction approach to derive the strong second-order sufficient condition from the Aubin property of the perturbed KKT system, in which the introduced lemma of alternative choices on cones plays an essential role. The result of this paper raises the question of whether the proposed approach can be applied to establish similar results for a broader range of (conic) optimization problems, in particular, for nonlinear semidefinite programming (SDP). This is especially relevant given that the strong regularity in nonlinear SDP is also equivalent to the strong second-order sufficient condition and constraint nondegeneracy established by Sun [29]. We leave this as one of our future research topics.

### v2

@@ 4 Equivalence of the Aubin property and the strong regularity
This section establishes the main result of this paper. The main step is to show that the strong second-order sufficient condition in Definition 4.1 holds. We will reformulate the Mordukhovich criterion for the Aubin property of ${\mathbb{S}}_{{\rm KKT}}$ to the necessary condition in Corollary 4.1. Moreover, by specifying a basis for the affine hull of the critical cone in Corollary 4.2, we can put our discussion in a space with a lower dimension (Lemma 4.3). Finally, by extensively utilizing Corollary 4.1 and Lemma 2.1, we obtain the strong second-order sufficient condition by Propositions 4.2 and 4.3, in which successive dimension reduction is involved.
Let ${\boldsymbol{x}}^{*}$ be a locally optimal solution to (1.1) with ${\boldsymbol{y}}^{*}=({\bm{\lambda}}^{*},{\bm{\mu}}^{*})$ being the associated multiplier. Note that any index $j\in\{1,\ldots,J\}$ belongs to one and only one of the following $6$ sets:
@@
For simplicity, $J_{1}\cup J_{2}$ is abbreviated to $J_{1,2}$, $J_{1}\cup J_{2}\cup J_{3}$ is abbreviated to $J_{1,2,3}$, etc. For an index set $I\subseteq\{1,2,3,4,5,6\}$, $|J_{I}|$ denotes the cardinality of the set $J_{I}$. Note that when $r_{j}=0$, $j$ must belong to $J_{1,2,6}$. Recall from [6, Lemma 25] that, for each $j=1,\ldots,J$, the tangent cone of ${\mathcal{Q}}_{j}$ at $g^{j}({\boldsymbol{x}}^{*})$ is given by
@@
Furthermore, by [6, Corollary 26] the critical cone ${\mathcal{C}}({\boldsymbol{x}}^{*})$ can be given by
@@
Moreover, according to [6, eq. (47)], the affine hull of the critical cone ${\mathcal{C}}({\boldsymbol{x}}^{*})$, i.e. the smallest subspace that contains the critical cone, can be formulated as
@@
In the following definition, the second-order and strong second-order sufficient conditions for (1.1) come from [6, Theorems 29 & 30].
@@
Let ${\boldsymbol{x}}^{*}$ be a stationary point of the nonlinear SOCP (1.1). We say that the second-order sufficient condition holds at ${\boldsymbol{x}}^{*}$ if there exists an associated multiplier $({\bm{\lambda}}^{*},{\bm{\mu}}^{*})$ such that
@@
where $\nabla_{{\boldsymbol{x}}{\boldsymbol{x}}}^{2}{\mathcal{L}}$ is the Hessian of ${\mathcal{L}}$ with respect to ${\boldsymbol{x}}$, and ${\mathcal{H}}({\boldsymbol{x}}^{*},{\bm{\mu}}^{*}):=\sum_{j=1}^{J}{\mathcal{H}}_{j}({\boldsymbol{x}}^{*},{\bm{\mu}}^{*})$ with
@@
We say that the strong second-order sufficient condition holds at ${\boldsymbol{x}}^{*}$ if there exists an associated multiplier $({\bm{\lambda}}^{*},{\bm{\mu}}^{*})$ such that
@@
Similarly, we say that the strong second-order necessary condition holds at ${\boldsymbol{x}}^{*}$ if there exists an associated multiplier $({\bm{\lambda}}^{*},{\bm{\mu}}^{*})$ such that
${\boldsymbol{d}},\left(\nabla^{2}_{{\boldsymbol{x}}{\boldsymbol{x}}}{\mathcal{L}}({\boldsymbol{x}}^{*},{\bm{\lambda}}^{*},{\bm{\mu}}^{*})+{\mathcal{H}}({\boldsymbol{x}}^{*},{\bm{\mu}}^{*})\right){\boldsymbol{d}}\right>\geq 0,\quad\forall\,{\boldsymbol{d}}\in{\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*})).$ (4.7)
Remark 4.1.
We will show that $\left(\nabla^{2}_{{\boldsymbol{x}}{\boldsymbol{x}}}{\mathcal{L}}({\boldsymbol{x}}^{*},{\bm{\lambda}}^{*},{\bm{\mu}}^{*})+{\mathcal{H}}({\boldsymbol{x}}^{*},{\bm{\mu}}^{*})\right){\boldsymbol{d}}\neq{\boldsymbol{0}}$ for all ${\boldsymbol{d}}\in{\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))\setminus\{{\boldsymbol{0}}\}$ (Corollary 4.1) if the Aubin property of ${\mathbb{S}}_{{\rm KKT}}$ holds. Consequently, we devote our main effort to proving that the strong second-order necessary condition (4.7) holds to obtain Proposition 4.3. In [3, Theorem 4.2], by assuming that the variational sufficiency for local optimality [31] holds, the authors established the equivalence between the Aubin property of ${\mathbb{S}}_{{\rm KKT}}$ and the strong regularity of (1.6) for stationary points. It can be observed from [31, Theorem 3 & Example 1] that for the conventional nonlinear programming, i.e., a special case of (1.2) with $r_{1}=\cdots=r_{J}=0$, the variational sufficiency for local optimality is a sufficient condition for the strong second-order necessary condition (4.7). Therefore, making such an assumption is very close to directly assuming the strong second-order sufficient condition (4.6). In contrast, we will derive the strong second-order sufficient condition solely from the Aubin property of ${\mathbb{S}}_{{\rm KKT}}$ without assuming the variational sufficiency as in [31].
@@ 4.1 Implications of the Aubin property
The following lemma, as an application of the reduction approach developed in [5, Section 3.4.4] for constraint systems, provides a convenient tool for characterizing the affine hull of the critical cone (4.4) via the null space of a given matrix.
@@
Suppose that Assumption 4.1 holds. Let $\Theta\in{\mathds{R}}^{n\times n}$ be the matrix given in (4.12), and $H\in{\mathds{R}}^{n\times\ell}$ be a matrix with full column rank such that ${\sf rge}H={\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$. Then the mapping ${{\mathbb{S}}}_{\rm GE}$ in (1.8) having the Aubin property at the origin for ${\boldsymbol{x}}^{*}$ implies that for all ${\bm{\nu}}\in{\mathds{R}}^{\ell}\setminus\{{\boldsymbol{0}}\}$ it holds that
@@ 4.2 Deriving the strong second-order sufficient condition
Since the solution mapping ${{\mathbb{S}}}_{\rm KKT}$ in (1.5) has the Aubin property at $({\boldsymbol{0}},{\boldsymbol{0}})$ for $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$, we know that the mapping ${\mathbb{S}}_{\rm GE}$ in (1.8) also has the Aubin property at the origin for ${\boldsymbol{x}}^{*}$. Then, by [10, Corollary 25] the constraint nondegeneracy condition (1.7) holds at ${\boldsymbol{x}}^{*}$ and the second-order sufficient condition holds (4.5) at ${\boldsymbol{x}}^{*}$. Therefore, Assumption 4.1 holds.
@@ 4.3 The equivalence theorem
Note that (1) is an immediate consequence of (2). On the other hand, if (1) holds, we know from [10, Corollary 25] that ${\boldsymbol{x}}^{*}$ is nondegenerate and the second-order sufficient condition (4.5) holds at ${\boldsymbol{x}}^{*}$. Consequently, it follows from Proposition 4.3 that the strong second-order sufficient condition (4.6) holds. Thus, by using [6, Theorem 30] we know that (2) holds. This completes the proof. ∎
Theorem 4.1 establishes the equivalence between the Aubin property and the strong regularity for the nonlinear SOCP (1.1) at local optimal solutions without requiring the strict complementarity. Moreover, Theorem 4.1 implies its counterpart for conventional nonlinear programming (((([12, Theorems 1, 4 &\&& 5])))) as a special case as one can take $r_{j}=0$ for all $j=1,\dots,J$.
According to Theorem 4.1, ${{\mathbb{S}}}_{\rm KKT}$ in (1.5) having the Aubin property at $({\boldsymbol{0}},{\boldsymbol{0}})$ for $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$ is also equivalent to many other conditions, such as (a) ${\boldsymbol{x}}^{*}$ is nondegenerate and the uniform second order growth condition ([5, Definition 5.16]) holds at ${\boldsymbol{x}}^{*}$ (by [6, Theorem 30]); (b) ${\boldsymbol{x}}^{*}$ is a nondegenerate and fully stable ([24, Definition 4.1]) locally optimal solution (by [24, Theorem 4.8]); and (c) Clarke’s generalized Jacobian of a KKT system is nonsingular (by [37, Theorem 3.1]). One may refer to the relevant references for details.
@@ 5 Conclusions
In this paper, we established the equivalence between the Aubin property of the perturbed KKT system and the strong regularity of the KKT system for nonlinear SOCP problems at locally optimal solutions. Our results extend beyond prior work by removing the restrictive condition $|J_{4}\cup J_{5}\cup J_{6}|\leq 1$ required in [27] and [26]. We achieved this by introducing a novel reduction approach to derive the strong second-order sufficient condition from the Aubin property of the perturbed KKT system, in which the lemma of alternative choices on cones we developed here plays an essential role. Our findings prompt further investigations into whether the approach introduced here could lead to similar equivalences for a broader range of (conic) optimization problems. Recently, for nonlinear semidefinite programming (SDP), [8] achieved the same equivalence after the announcement of this work, but their approach leverages properties specific to nonlinear SDP. While SOCPs can be reformulated as SDPs, simultaneous satisfaction of constraint nondegeneracy can fail [34, 5], implying that the results in [8] cannot carry over to SOCPs111 To see this, one may consider the nonlinear SOCP of minimizing $f(x_{0},x_{1}):=\frac{1}{2}(x_{0}^{2}+2x_{0}+x^{2}_{1})$ subject to $(x_{0},x_{1})\in\{(x_{0},x_{1})\mid x_{0}\geq|x_{1}|\}$, together with its equivalent SDP reformulation of minimizing $f(x_{0},x_{1})$ subject to $\begin{pmatrix}x_{0}&x_{1}\\ x_{1}&x_{0}\end{pmatrix}\succeq 0$, as an illustrative example.. The equivalence between the Aubin property and the strong regularity for general optimization problems, particularly those with non-polyhedral $C^{2}$-cone reducible constraints, remains an open question. We leave this as one of our future research topics.